Reports 1-1 of 1 Clear search Modify search
Virgo Runs (O3)
berni - 15:01 Monday 13 May 2019 (45875) Print this report
operator report
ITF found in Science mode.
At 6:59 UTC the ITF unlocked because of the high wind.
At 7:10 UTC I enabled the GIPC configuration for high wind condition.
At 7:37 UTC the ITF was back in Science mode but it unlocked at 7:57 still because high wind.

After this unlock the wind was very high and prevented to achieve Science mode; Paolo is working trying improve the performance in this condition.
We experienced also some unlocks of the squeezer which prevented to stay in LOW_NOISE_3_SQZ.

At the moment the best configuration we could reach is LOW_NOISE_3 locked with input mirrors and GIPC for high wind enabled; Hrec_Range_BNS around 43 Mpc.
Due to the not standard configuration the ITF has been left in the default mode LOCKED, (see attached plot).

ISC
BS AA full bandwidth restored, see entry BS loop back in full bandwidth; no more pending action.
Images attached to this report
Comments to this report:
ruggi - 16:12 Monday 13 May 2019 (45877) Print this report

To be precise, BS loop full bandwidth is still under observation: the engagement has been done manually, the automatic engagement at LN3 is still commented.

ruggi - 17:14 Monday 13 May 2019 (45879) Print this report

BS full bandwidth is back in operation also in the automation.

Search Help
×

Warning

×